OpenStudy (anonymous):

If the sin 60° is square root 3 over 2, then the cos ____° = _____. 30° ; 1 over 2 60° ; square root 2 over 2 30° ; square root 3 over 2 60° ; 1

OpenStudy (anonymous):

======================================... tan(x) = sin(x)/cos(x) Therefore: tan(60) = sin(60)/cos(60) = (√3/2)/(1/2) = √3 ======================================... sec(x) = 1/cos(x) Therefore: sec(60) = 1/(cos(60) = 1/(1/2) = 2 ======================================... cot(x) = cos(x)/sin(x) Therefore: cot(60) = cos(60)/sin(60) = (1/2)/(√3/2) = 1/√3 ======================================...